OpenStudy (hkmiu):

How do you write a recursive formula for a sequence whose first three terms are 1, 1 and 3? I feel like the question didn't give enough numbers.

OpenStudy (anonymous):

hey hottie

OpenStudy (ybarrap):

This is one pattern that I can think of, which might work 1 +1 =2 1+2=3 3+3=9 9+4=12 12+5=17 ...

ganeshie8 (ganeshie8):

infinite patterns are possible as there is not sufficient info. another pattern is :- \(\large a_1 = 1\) \(\large a_{n+1} = 3^{n-1}\)

OpenStudy (ybarrap):

yes, that's true also

OpenStudy (hkmiu):

@ybarrap that was what I was thinking too. But there wasn't any 2's in the pattern. If there was it bould be a fibonacci sequence. idk. @ganeshie8, for the formula, for a2 I got 3. But 3 was supposed to be a3.. It's a bit confusing.

ganeshie8 (ganeshie8):

\(\large a_1 = 1\) \(\large a_2 = a_{1+1} = 3^{1-1} = 3^0 = 1\) \(\large a_3 = a_{2+1} = 3^{2-1} = 3^1 = 3\)

OpenStudy (hkmiu):

That made sense when you explained it! Thanks again! @ganeshie8

ganeshie8 (ganeshie8):

np :) but you're right, there is not sufficient data here, if we think a bit we can form many recursive formulas like above

ganeshie8 (ganeshie8):

another one :- \(\large a_1 = 1\) \(\large a_{n+1} = 2^n-1\)

OpenStudy (hkmiu):

Aaaaaaah, :(. I don't like recursion and special sequences! I've been stuck on this lesson for a good 3 hours now.

OpenStudy (hkmiu):

I submitted my work & got a 90. It gave me the possible answer ----> http://media.education2020.com/evresources//423123.jpg It doesn't look the same to me but I got it right. (or maybe my eyes and brain are just too tired to realize)...

ganeshie8 (ganeshie8):

congrats ! :) yes that works too...